A 0.004 mol dm-3 solution of epsom salts contains 1.072g of the salt on 1dm3 of solution. Find the mass of 1 mol of epsom salts. If the formula of epsom salt is represented by MSO4•7H2O, find the relative atomic mass of M.
1
Expert's answer
2018-11-27T07:06:05-0500

Answer on question 83317

A 0.004 mol dm-3 solution of epsom salts contains 1.072g of the salt on 1dm3 of solution. Find the mass of 1 mol of epsom salts. If the formula of epsom salt is represented by MSO4•7H2O, find the relative atomic mass of M.

Solution

a) Find the mass of 1 mol of epsom salts:

m = M*n, M = m/n

M(MSO4*7H2O) = 1.072 g/ 0.004 mol = 268 g/mol

b) Mr(MSO4*H2O) = Ar(M) + 32.07 + 16.00*4 + 7*(2*1.01 + 16.00) = Ar(M) + 222.21

Mr = M

Ar(M) +222.21 = 268

Ar(M) = 45.79

Answer: a) 268 g/mol; b) 45.79